# Transverse beam instabilities in low-emittance booster synchrotrons

### v1

@@ III.2 Simulation results
The simulations were done at the multibunch nominal charge 3 nC and at the maximum technologically achievable charge for SOLEIL II booster of 10 nC. These charges are distributed over 104 bunches of the total length of 295 ns, corresponding to the injection scheme for the 416-bunch operation mode of the storage ring. Knowing that the booster harmonic number is 184, this filling pattern leaves 80 empty buckets after the bunch train.
@@
On the contrary, in the case where the ADTS is included shown in Fig. 10(b), the beam does not suffer from instability, whether at 3 or 10 nC and regardless of chromaticity. It can then be inferred that the beam is also stable for other charges below 10 nC. This favorable outcome is due to the large tune spread induced by the ADTS [21]. It emphasizes the beneficial side of the nonlinearity that comes with a 4GLS booster. At the same time, it reminds us that adding nonlinearity in tracking simulation for such a machine is important due to the large impact it can have on the results.
@@ III.3 Head-tail damping
The explanation does involve the single-bunch scale mechanism usually called the head-tail damping. It refers to the situation where the TCBI is suppressed by shifting the chromaticity to a positive value. It does not compete with instability with its own damping rate like, for example, the synchrotron radiation, but rather directly reduces the instability growth rate. This is because the theoretical TCBI growth rate is proportional to effective impedance [30], which is a sum of the transverse impedance weighted by the bunch power spectrum resulting from the head-tail oscillation. In other words, it is the net quantity of the impedance that can be sampled by the bunch frequency. One can see it as an overlapping area between the impedance and the bunch spectrum. The impedance’s shape, the head-tail oscillation mode, and the chromaticity can have an impact on how these two variables overlap.
@@
In case of the SOLEIL II booster, its vertical dipolar impedance is shown in Fig. 12 and Table 3 summarizes the effective impedance and the theoretical TCBI growth rate at the injection energy 150 MeV at chromaticities 1–3 for the first three head-tail modes (see the equations used for the calculation in Appendix B). For chromaticities 2 and 3, as we have deducted from Fig. 11, we are certain that the dominant mode that represents the instability is , in agreement with this calculation. For chromaticity 1, however, we cannot know for sure which mode dominates the beam since the instability did not occur. In any case, if we pick its strongest mode and compare it to the same mode of the other chromaticities, we can see that the growth rate at chromaticity 1 is indeed smaller than that at chromaticity 2 and 3.
@@
1 | 0 | -3419 | |
1 | 437 | ||
2 | 328 | ||
2 | 0 | -5233 | |
1 | 704 | ||
2 | 551 | ||
3 | 0 | -6901 | |
1 | 834 | ||
2 | 755 |
This is effectively due to the shape of the impedance. In a 4GLS machine, one might expect the resistive-wall impedance to dominate since the vacuum chamber radius tends to be small and use only this to estimate the machine’s impedance. But, Fig. 12 has shown us that our booster is actually composed of two main parts; a resistive-wall impedance which peaks at zero frequency and a broad-band impedance which peaks around 450–500 MHz. The first part, of course, comes from the vacuum chamber around the machine. The second part appears due to the presence of three ceramic chambers whose inner wall is coated with a 200 nm thick titanium layer, used to accommodate the fast kicker magnets at the injection and extraction points [10].
Such an impedance shape damps the head-tail mode , whose power amplitude is relatively large, and gives rise to other higher-order modes with a much lower amplitude but have a better overlap with the impedance. Without the broad-band part in the impedance model, the effect of head-tail damping will be much weaker so that it cannot stabilize the beam. This is what we can observe in Fig. 13 which shows the emittance along the ramp of a 3 nC beam without ADTS and without broad-band impedance, only a resistive-wall impedance. When compared to Fig. 10(a), we can notice a larger emittance at zero chromaticity and that the beam at chromaticity 1 is not anymore stable.
In summary, the results in this section tell us that broad-band impedance can be beneficial for TCBI since it can reduce the effective impedance by damping the powerful head-tail mode 0 and exciting weaker higher-order modes. The particular shape of the higher-mode bunch spectrum can then result in a nonlinear relationship between the chromaticity and TCBI growth rate. This, however, does not mean that broad-band resonators should be imprudently introduced in the machine since they can worsen the instabilities in the single-bunch regime, such as increasing TMCI and head-tail growth rate or reducing microwave instability threshold. Instead, it tells us that care should be taken when one wants to model a machine’s impedance for a TCBI study, not to neglect possible broad-band impedance sources.
@@ IV Conclusion
For the multibunch regime, the studies have shown that ADTS can be an effective mechanism against TCBI as it can sustain the beam’s stability throughout the ramp in the SOLEIL II case because of Landau damping similar to the single-bunch case.
Lastly, we have shown that a broad-band impedance at high frequency can result in a suppression of TCBI at slightly positive chromaticity due to the head-tail damping effect.

### v2

@@ III.2 Simulation results
The simulations were carried out at the nominal multi-bunch charge 3 nC (equivalent to 5.7 mA) and at the maximum technologically achievable charge for the SOLEIL II booster of 10 nC (19.2 mA). These charges are distributed over 104 bunches of the total length of 295 ns, corresponding to the injection scheme for the 416-bunch operation mode of the storage ring. Knowing that the booster harmonic number is 184, this filling pattern leaves 80 empty buckets after the bunch train.
The cavity voltage was modeled as a sinusoidal wave without taking into account the transient beam loading effect since the beam current is small, even at the maximum charge.
@@
In contrast, in the case where ADTS is included as shown in Fig. 10(b), the beam does not suffer from instability, whether at 3 or 10 nC and regardless of chromaticity. It can then be inferred that the beam is also stable for other charges below 10 nC. This favorable outcome is due to the large tune spread induced by the ADTS [23]. It emphasizes the beneficial side of the nonlinearity that comes with a 4GLS booster. At the same time, it reminds us that adding nonlinearity in tracking simulation for such a machine is important due to the large impact it can have on the results.
@@ III.3 Head-tail damping
The explanation does involve the single-bunch scale mechanism usually called the head-tail damping. It refers to the situation where the TCBI is suppressed by shifting the chromaticity to a positive value. It does not compete with instability with its own damping rate like, for example, the synchrotron radiation, but rather directly reduces the instability growth rate. This is because the theoretical TCBI growth rate is proportional to effective impedance [37], which is a sum of the transverse impedance weighted by the bunch power spectrum resulting from the head-tail oscillation. In other words, it is the net quantity of the impedance that can be sampled by the bunch frequency. One can see it as an overlapping area between the impedance and the bunch spectrum. The shape of the impedance, the head-tail oscillation mode, and the chromaticity can have an impact on how these two variables overlap.
@@
In case of the SOLEIL II booster, its vertical dipolar impedance is shown in Fig. 12 and Table 3 summarizes the effective impedance and the theoretical TCBI growth rate at the injection energy 150 MeV at chromaticities 1–3 for the first three head-tail modes (see the equations used for the calculation in Appendix B). For chromaticities 2 and 3, as we have deduced from Fig. 11, the dominant mode that represents the instability is , in agreement with this calculation. For chromaticity 1, however, we cannot know for sure which mode dominates the beam since the instability does not occur. In any case, if we pick its strongest mode and compare it with the same mode of the other chromaticities, we can see that the growth rate at chromaticity 1 is indeed smaller than that at chromaticity 2 and 3.
@@
| 1 | 0 | -3419 | |
| 1 | 437 | ||
| 2 | 328 | ||
| 2 | 0 | -5233 | |
| 1 | 704 | ||
| 2 | 551 | ||
| 3 | 0 | -6901 | |
| 1 | 834 | ||
| 2 | 755 |
This is effectively due to the shape of the impedance. In a 4GLS machine, one might expect the resistive-wall impedance to dominate since the vacuum chamber radius tends to be small and use only this to estimate the machine impedance. But, Fig. 12 has shown us that our booster is actually composed of two main parts; a resistive-wall impedance which peaks at zero frequency and a broad-band impedance which peaks around 450–500 MHz. The first part, of course, comes from the vacuum chamber around the machine. The second part appears due to the presence of three ceramic chambers whose inner wall is coated with a 200 nm thick titanium layer, used to accommodate the fast kicker magnets at the injection and extraction points [10].
Such an impedance shape damps the head-tail mode , whose power amplitude is relatively large, and gives rise to other higher-order modes with a much lower amplitude but have a better overlap with the impedance. Without the broad-band part in the impedance model, the effect of head-tail damping will be much weaker so that it cannot stabilize the beam. This is what we can observe in Fig. 13 which shows the emittance along the ramp of a 3 nC beam without ADTS and without the broad-band impedance, only the resistive-wall one. When compared to Fig. 10(a), we can notice a larger emittance at zero chromaticity and that the beam at chromaticity 1 is not anymore stable.
In summary, the results in this section tell us that broad-band impedance can be beneficial for TCBI since it can reduce the effective impedance by damping the powerful head-tail mode 0 and exciting weaker higher-order modes. The particular shape of the higher-mode bunch spectrum can then result in a nonlinear relationship between the chromaticity and TCBI growth rate. This, however, does not mean that broad-band resonators should be imprudently introduced in the machine since they can worsen the instabilities in the single-bunch regime, such as increasing TMCI and head-tail growth rate or reducing microwave instability threshold. Instead, it tells us that care should be taken when one wants to model a machine impedance for a TCBI study, not to neglect possible broad-band sources.
@@ IV Conclusion
For the multibunch regime, the studies have shown that ADTS can be an effective mechanism against TCBI as it can sustain the beam stability throughout the ramp in the SOLEIL II case because of Landau damping similar to the single-bunch case.
Lastly, we have confirmed that a broad-band impedance at high frequency can result in a suppression of TCBI at slightly positive chromaticity due to the head-tail damping effect, as explained in [37].